OpenStudy (anonymous):

Sketch the plane curve defined by the given parametric equations and find a corresponding x-y equation for the curve. x=3+8t, y=7t. Answer should be formatted: y=___ x - ___

OpenStudy (campbell_st):

make t the subject in both equations \[t = \frac{x - 3}{8}... and ... t = \frac{y}{7}\] substituting you get \[\frac{y}{7} = \frac{x - 3}{8}\] or \[y = \frac{7(x - 3)}{8}\] just simplify